OpenStudy (anonymous):

what is the equation of the line that is perpendicular to y=2/3x -6 and passes through the point (6,-3)?

OpenStudy (anonymous):

slope is \[-\frac{3}{2}\] point is (6,-3) use \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

to get \[y-(-3)=\frac{3}{2}(x-6)\]

OpenStudy (anonymous):

multiply out to get \[y+3=-\frac{3}{2}x+9\] or \[y=-\frac{3}{2}x+6\]

OpenStudy (anonymous):

wait so the final answer is wut?